# Measuring entanglement without local addressing in quantum many-body simulators via spiral quantum state tomography

### v2

@@ Results
The spin-1/2 Heisenberg model, a foundational model for studying quantum magnetism, has been realized in the strong interaction regime of systems consisting of two-component ultracold Fermi Boll2016-wd ; Mazurenko2017-yb ; Brown2017-fp or Bose Fukuhara2015-sn ; Jepsen2020-ze ; Jepsen2021-xl ; Sun2021-mv atoms trapped in an optical lattice. Antiferromagnetic correlations have been observed in both one Boll2016-wd and two Cheuk2016-mw ; Mazurenko2017-yb ; Brown2017-fp dimensions using spin-selective imaging with a quantum-gas microscope (QGM) Gross2021-iv , which enables site-resolved longitudinal ($Z$-basis) measurements. The correlation length of the observed antiferromagnetic correlations extends to approximately 10 sites in Ref. Mazurenko2017-yb . Transverse ($X$- or $Y$-basis) measurements are also possible by applying a global $\pi/2$ rotation before imaging Mazurenko2017-yb ; Brown2017-fp . In these typical optical-lattice systems, however, the complete reconstruction of density matrices, which conventionally requires site-resolved rotations of spin directions, has not yet been achieved, partly due to the short lattice spacing imposed by the laser wavelength used to form the optical lattice.
@@
Figure 3: Magnetic field gradient method for spiral QST. a Magnetic field gradient used to create spiral-spin settings. Typical experimental noise arises from fluctuations in the strength of the magnetic field over time. These fluctuations cause variations in the zero point of the gradient, modeled as Gaussian fluctuations with a mean of 0 and a standard deviation $\sigma_{\rm zp}$ (in units of lattice spacing). b Protocol for numerical simulations of spiral QST with magnetic field gradient. We simulate a realistic experimental process to measure observables by averaging the outcomes over many shots of repeatedly prepared target states for each experimental setting of spiral plane $\alpha\beta=XY$, $YZ$, or $ZX$ and pitch $q$. Each repetition is subject to fluctuations in the magnetic field gradient, introducing noise into the expectation values of the spiral operators.
@@
Below, we numerically simulate the spiral compressed-sensing QST for Heisenberg spin systems, assuming that spiral measurements are implemented through the combined application of a magnetic-field gradient $H_{\rm grad}$ as in Eq.(3Hild2014-mz ; Jepsen2020-ze ; Jepsen2021-xl and global $\pi/2$ Rabi pulses Fukuhara2015-sn ; Mazurenko2017-yb ; Brown2017-fp , followed by $Z$-basis QGM measurements. To account for potential experimental errors beyond the abstract noise models parameterized by $\gamma$ and $\sigma$ in the previous subsection, we consider more concrete scenarios. In realistic experimental settings, a significant source of measurement noise could arise from fluctuations in the zero point of the magnetic field gradient, modeled as $-B(i-i_{0})\rightarrow-B(i-i_{0}-\delta_{\rm zp})$, where $\delta_{\rm zp}$ represents the zero-point displacement in units of the lattice spacing (see Fig. 3a). This displacement leads to an unwanted phase shift in the spiral structures, causing the experimentally obtained values to correspond to spiral operators with $q(i-i_{0})$ [Eq. (2)] replaced by $q(i-i_{0}-\delta_{\rm zp})$. We assume that $\delta_{\rm zp}$ is Gaussian distributed with a mean of 0 and a standard deviation $\sigma_{\rm zp}$. It should also be noted that the accurate determination of the expectation values requires a sufficiently large number of repetitions of the same measurements for each spiral setting. For each experimental setup — defined by a specific spin-spiral plane and pitch angle $q$ — we simulate multiple QGM shots, each affected by the zero-point fluctuations of the magnetic field gradient as described above. An adequate number of repetitions is necessary to capture the quantum nature of the state, namely the quantum fluctuations inherent in measuring each spiral operator.
Figure 4: Spiral QST for the ground state of the NN Heisenberg chain. Sample-averaged fidelity and trace distance for spiral compressed-sensing QST of the ground state of the 8-site Heisenberg chain with NN antiferromagnetic spin-exchange coupling, plotted as functions of the number of measurements $m$ scaled by $d^{2}=4^{8}$. In (a), the results for different amplitudes of magnetic field fluctuations ($\sigma_{\rm zp}=0.01,0.1,0.5$) are compared, with the number of repetitions for evaluating each expectation value fixed at ${\rm reps}=500$. The dependence on $\sigma_{\rm zp}$ is essentially absent, which is supported by the symmetry argument given in the main text. In (b), the results for different numbers of repetitions (${\rm reps}=100,500,1000$) are compared (the amplitude of magnetic field fluctuations is chosen as $\sigma_{\rm zp}=0.1$ for concreteness). We adopt the strategy of progressively including expected relevant pitch angles; the data points with the fewest measurements correspond to the spiral QST with $l=0$, followed by $l=0,1$, $l=0,1,2$, and so on.
@@
First, we consider the test case of the ground state $|\Psi_{0}\rangle$ of the $N=8$ Heisenberg model (8) with only nearest-neighbor (NN) antiferromagnetic coupling ($J_{1}=1$ and $J_{p\neq 1}=0$), which can easily be computed using the exact diagonalization method. For highly symmetric states, such as the ground state of typical condensed-matter Hamiltonians, the majority of the QST weight is expected to concentrate on spiral operators associated with a limited number of pitch angles. This observation suggests a natural strategy: constructing an expansion that prioritizes the most relevant pitch angles and planning the experiment based on this hierarchy. Accordingly, we perform spiral QST by progressively including pitch angles from the set defined in Eq. (4), starting with ${l=0}$, followed by ${l=0,1}$, ${l=0,1,2}$, and so on (see Methods for an analysis of relevant pitch angles).
@@
In Fig. 4a, we present the fidelity and the trace distance between the target state $\rho=|\Psi_{0}\rangle\langle\Psi_{0}|$ and the reconstructed state $\rho^{\rm rec}$ obtained via the spiral QST, with the number of repetitions (denoted by “reps”) fixed at ${\rm reps}=500$ for evaluating each expectation value. We compare the results for different amplitudes of magnetic field fluctuations ($\sigma_{\rm zp}=0.01,0.1,0.5$); other sources of noise are absent. First, we observe that the dependence on $\sigma_{\rm zp}$ appears to be absent, except for variations attributable only to statistical errors. Second, our “relevant-pitch-angles” strategy appears to be highly effective. The fidelity reaches approximately $0.98$ for $m/d^{2}\approx 0.1$, which corresponds to only 20-30 independent measurement setups, considering that one can obtain $d=2^{N}$ independent correlation functions from each experimental setup. The data points with the fewest measurements correspond to the results obtained by using just $m=3\times 2^{8}-2$ expectation values, based merely on three experimental setups of uniform $X$-, $Y$-, and $Z$-measurements (i.e., $l=0$ or $q=0$) in the SVT method. This condition already achieves very good fidelity and trace distance, likely due to the simplicity of the ground state of the NN Heisenberg model. Interestingly, the fidelity decreases (and the trace distance increases) when additional measurements are included in the small $m/d^{2}$ range; this may simply be due to the possibility that a better optimization of the SVT algorithm for $l=0$ was achieved in our attempts, but further clarifications would be worthwhile.
@@
In Fig. 4b, we compare the results obtained by varying the number of repetitions for each experimental setting (i.e., the number of QGM shots after each unitary transformation corresponding to spiral operators defined by a specific pitch angle $q$ and spin plane) while keeping the amplitude of magnetic field fluctuations fixed at $\sigma_{\rm zp}=0.1$, although a different value would not change the results, as discussed above. We observe that this dependence is significant, at least below a certain threshold, although this sensitivity is a general issue for QST. Our results suggest that ${\rm reps}\gtrsim 500$ is required to achieve convergence.
@@
In Fig. 5, we present the results for the ground state of the $N=8$ Heisenberg + DM chain with $D=-J_{1}=1$ and $J_{p\neq 1}=0$, adopting the same sequence of pitch angles as in the previous case without DM interactions. As seen in the figure, uniform measurements with $l=0$ alone are insufficient for accurately reconstructing the ground state in this case. The DM interaction introduces additional complexities, requiring the inclusion of finite pitch angles to capture the non-collinear spin correlations present in the ground state. Specifically, we identified the most relevant pitch angle (taken individually) in this case as $l=3$ or $q=\pi/2$ (see Methods).
For larger-size system realized in experiments, convergence might be achieved with fewer measurements by employing a strategy that prioritizes the most relevant pitch angles, guided by simulations on smaller systems. Nevertheless, as demonstrated in Fig. 5, the compressed sensing with the SVT method already performs remarkably well, even when the most relevant pitch angle ($l=3$ in this case) is not explicitly included.
In this case, as expected from the absence of the global SU(2) symmetry, an appreciable dependence on $\sigma_{\rm zp}$ is indeed present, although the results converge very well around and below $\sigma_{\rm zp}=0.1$. The fluctuations with $\sigma_{\rm zp}=0.01$ roughly correspond to an experimental accuracy of 10 µGauss, which, while attainable, remains challenging. These results suggest that such extremely precise control is not necessary for our protocol, at least under the assumption of a symmetric zero-point distribution (such as a Gaussian).
Detection of entanglement measures
@@
for the reduced density matrix $\rho_{A}\equiv{\rm Tr}_{\bar{A}}(\rho)$ of subsystem $A$, when the whole system is in a pure quantum state. Here, ${\rm Tr}_{\bar{A}}(\cdot)$ denotes the partial trace with respect to the complement of $A$. The experimental detection of the entanglement entropy is of paramount importance, as it provides a versatile and precise measure of quantum correlations across various physical contexts. In condensed-matter physics, it is instrumental in identifying quantum critical points Osterloh2002-ng ; Eisert2010-ko and topological order Kitaev2006-vm ; Wen2017-ax . In quantum field theory, it reveals underlying structural properties, including the central charge and operator content, through its relationship with the entanglement spectrum Calabrese2004-br ; Lauchli2013-wz . Moreover, entanglement entropy is deeply connected to black hole entropy in gravitational theory Kabat1995-bs ; Solodukhin2011-rv and serves as a crucial indicator of information spread and thermalization processes in quantum thermodynamics Kaufman2016-db ; DAlessio2016-zf .
@@
Despite their critical importance, the experimental determination of entanglement measures remains challenging, particularly in large-scale systems. Direct measurements of entanglement have been rare, especially in more extensive systems. For instance, in cold-atom systems, while the second Rényi entanglement entropy $S_{2}(\rho_{A})$ has been measured by letting two identical copies of the quantum state in a Bose–Hubbard system interfere, such measurements have so far been implemented in systems with a modest number of sites Islam2015-rn ; Kaufman2016-db .
@@
Figure 6: Spiral QST for the reduced density matrix. a Illustration of the bipartition of a system with $N$ sites into subsystem $A$, consisting of $N_{A}$ sites, and its complement $\bar{A}$. b Sample-averaged fidelity and trace distance for spiral compressed-sensing QST of the reduced density matrix $\rho_{A}={\rm Tr}_{\bar{A}}(\rho)$, where $N_{A}=7$, plotted as functions of the number of measurements $m$ scaled by $d^{2}=4^{7}$. The full density matrix $\rho$ corresponds to the ground state of the 14-site Heisenberg chain with NN antiferromagnetic spin-exchange coupling. The number of repetitions for evaluating each expectation value is set to ${\rm reps}=500$. The data points with the fewest measurements correspond to the spiral QST with $l=0$, followed by $l=0,1$, $l=0,1,2$, and so on.
@@
We now demonstrate how entanglement measures in a quantum state can be extracted using spiral QST, which has the potential to overcome the challenge. Notably, to obtain the reduced density matrix $\rho_{A}$ via QST, it is not necessary to first reconstruct the entire pure state and then take a partial trace. Instead, the QST process can be applied directly to the measurement outcomes from the targeted subsystem $A$. To illustrate, we perform numerical simulations as follows: We consider a spin chain of length $N=14$, assumed to be prepared in the ground state $|\Psi_{0}\rangle$ of the Heisenberg model (8), discussed in the previous subsection. Focusing on a subsystem $A$ consisting of sites form $i=1$ to $i=N_{A}$ [see Fig. 6a], we simulate the reconstruction of the reduced density matrix $\rho_{A}^{\rm rec}$ on $A$ via the spiral compressed-sensing QST. The von Neumann and Rényi entanglement entropies are then calculated from the reconstructed $\rho_{A}^{\rm rec}$, and compared with their theoretical values.
@@
A priori, it is not clear whether spiral QST (or, more generally, compressed sensing) is suitable for a reduced density matrix, as it typically represents a mixed state of generic rank. Therefore, we begin by comparing $\rho_{A}^{\rm rec}$ with the theoretical reduced density matrix $\rho_{A}$, which is obtained by tracing out the complement of $A$ from $\rho=|\Psi_{0}\rangle\langle\Psi_{0}|$. In Fig. 6b, we present the fidelity and trace distance between these matrices, with the subsystem size set to $N_{A}=7$ (half of the entire system). We set the number of repetitions for expectation value calculations to ${\rm reps}=500$. The value of $\sigma_{\rm zp}$ does not affect the results because the reduced density matrix retains the SU(2) symmetry of the ground state $|\Psi_{0}\rangle$ of the entire system. The results demonstrate that spiral QST is indeed effective also for the reduced density matrix, exhibiting a similar behavior to the case shown in Fig. 4. This can be attributed to the strong hierarchy of eigenvalues, where a few dominant ones significantly outweigh the others, a characteristic commonly observed in physical systems, which renders this reduced density matrix effectively low-rank.
@@
Now let us move on to the entanglement measures given in Eqs.(10) and (11). In Figs. 7a-c, we present the sublattice-size $N_{A}$ dependence of $S_{\rm vN}(\rho_{A})$, $S_{2}(\rho_{A})$, and $S_{4}(\rho_{A})$, respectively, obtained from the reconstructed reduced density matrix $\rho_{A}^{\rm rec}$, for the case of ${\rm reps}=500$ (for concreteness $\sigma_{\rm zp}=0.1$ is chosen), alongside the theoretical values. The target state describes a subsystem of size $N_{A}$ of the 14-site NN antiferromagnetic Heisenberg chain ($J_{1}=1$ and $J_{p\neq 1}=0$) in its ground state. For the reconstruction, we used only the measurements with $l=0$ ($m=3\times 2^{N_{A}}-2$), as the fidelity and trace distance are already sufficiently high and low, respectively, as seen in Fig. 6b.
@@
Moving beyond the previous simple model, we now demonstrate a more interesting case with antiferromagnetic NN coupling $J_{1}>0$ and next-nearest-neighbor (NNN) coupling $J_{2}>0$. These interactions compete with each other, leading to frustrations among the spins, which is expected to induce a larger entanglement in the ground state. In experiments with optical lattices, such a model can be realized by arranging the laser beams in a way that creates a zig-zag chain configuration Zhang2015-bh . Note that the zig-zag structure does not interfere with the procedure involving a magnetic field gradient along the chain direction.
We present the results for $J_{1}=J_{2}=1$ and $J_{p\neq 1,2}=0$ in Figs. 8a-c. Although the von Neumann entropy still proves to be relatively challenging due to the logarithmic weighting that amplifies the errors, these plots reveal how well the spiral QST can capture the entanglement properties of the ground state of quantum many-body systems. These demonstrations provide valuable insights into the effectiveness of the method in important and interesting scenarios, such as frustrated quantum systems.
Discussion

### v3

@@ III.1 Spiral-QST for typical ground states
Figure 3: Magnetic field gradient method for spiral QST. (a) Magnetic field gradient used to create spiral-spin settings. Typical experimental noise arises from fluctuations in the strength of the magnetic field over time. These fluctuations cause variations in the zero point of the gradient, modeled as Gaussian fluctuations with a mean of 0 and a standard deviation $\sigma_{\rm zp}$ (in units of lattice spacing). (b) Protocol for numerical simulations of spiral QST with magnetic field gradient. We simulate a realistic experimental process to measure observables by averaging the outcomes over many shots of repeatedly prepared target states for each experimental setting of spiral plane $\alpha\beta=XY$, $YZ$, or ZXZX and pitch $q$. Each repetition is subject to fluctuations in the magnetic field gradient, introducing noise into the expectation values of the spiral operators.
The spin-1/2 Heisenberg model, a foundational model for studying quantum magnetism, has been realized in the strong interaction regime of systems consisting of two-component ultracold Fermi [47, 35, 36] or Bose [34, 40, 41, 48] atoms trapped in an optical lattice. Antiferromagnetic correlations have been observed in both one [47] and two [49, 35, 36] dimensions using spin-selective imaging with a quantum-gas microscope (QGM) [50], which enables site-resolved longitudinal ($Z$-basis) measurements. The correlation length of the observed antiferromagnetic correlations extends to approximately 10 sites in Ref. [35]. Transverse ($X$- or $Y$-basis) measurements are also possible by applying a global $\pi/2$ rotation before imaging [35, 36]. In these typical optical-lattice systems, however, the complete reconstruction of density matrices, which conventionally requires site-resolved rotations of spin directions, has not yet been achieved, partly due to the short lattice spacing imposed by the laser wavelength used to form the optical lattice.
@@
Below, we numerically simulate the spiral compressed-sensing QST for Heisenberg spin systems, assuming that spiral measurements are implemented through the combined application of a magnetic-field gradient $H_{\rm grad}$ as in Eq.(3[39, 40, 41, 42] and global $\pi/2$ Rabi pulses [34, 35, 36], followed by $Z$-basis QGM measurements. To account for potential experimental errors beyond the abstract noise models parameterized by $\gamma$ and $\sigma$ in the previous subsection, we consider more concrete scenarios. In realistic experimental settings, a significant source of measurement noise could arise from fluctuations in the zero point of the magnetic field gradient, modeled as $B(i-i_{0})\rightarrow-B(i-i_{0}-\delta_{\rm zp})$, where $\delta_{\rm zp}$ represents the zero-point displacement in units of the lattice spacing [see Fig. 3(a)]. This displacement leads to an unwanted phase shift in the spiral structures, causing the experimentally obtained values to correspond to spiral operators with $q(i-i_{0})$ [Eq. (2)] replaced by $q(i-i_{0}-\delta_{\rm zp})$. We assume that $\delta_{\rm zp}$ is Gaussian distributed with a mean of 0 and a standard deviation $\sigma_{\rm zp}$. It should also be noted that the accurate determination of the expectation values requires a sufficiently large number of repetitions of the same measurements for each spiral setting. For each experimental setup — defined by a specific spin-spiral plane and pitch angle $q$ — we simulate multiple QGM shots, each affected by the zero-point fluctuations of the magnetic field gradient as described above [see Fig. 3(b)]. An adequate number of repetitions is necessary to capture the quantum nature of the state, namely the quantum fluctuations inherent in measuring each spiral operator.
@@
First, we consider the test case of the ground state $|\Psi_{0}\rangle$ of the $N=8$ Heisenberg model (8) with only nearest-neighbor (NN) antiferromagnetic coupling ($J_{1}=1$ and $J_{p\neq 1}=0$), which can easily be computed using the exact diagonalization method. For highly symmetric states, such as the ground state of typical condensed-matter Hamiltonians, the majority of the QST weight is expected to concentrate on spiral operators associated with a limited number of pitch angles; in addition, the pitch angles are related by the Hamiltonian’s symmetry (which is not spontaneously broken at finite size). This observation suggests a natural strategy: constructing an expansion that prioritizes the most relevant pitch angles as well as the symmetries, and planning the experiment based on this hierarchy. Accordingly, we perform spiral QST by progressively including pitch angles from the set defined in Eq. (4), starting with $\{l=0\}$, followed by $\{l=0,1,-1\}$, $\{l=0,1,-1,2,-2\}$, and so on (see Appendix B for an analysis of relevant pitch angles). Note that all these sets respect inversion symmetry.
Figure 4: Spiral QST for the ground state of the NN Heisenberg chain. Sample-averaged fidelity and trace distance for spiral compressed-sensing QST of the ground state of the 8-site Heisenberg chain with NN antiferromagnetic spin-exchange coupling, plotted as functions of the number of measurements $m$ scaled by $d^{2}=4^{8}$. In (a), the results for different amplitudes of magnetic field fluctuations ($\sigma_{\rm zp}=0.01,0.1,0.5$) are compared, with the number of repetitions for evaluating each expectation value fixed at ${\rm reps}=500$. The dependence on $\sigma_{\rm zp}$ is essentially absent, which is supported by the symmetry argument given in the main text. In (b), the results for different numbers of repetitions (${\rm reps}=100,500,1000$) are compared (the amplitude of magnetic field fluctuations is chosen as $\sigma_{\rm zp}=0.1$ for concreteness). We adopt the strategy of progressively including expected relevant pitch angles in symmetric fashion; the data points with the fewest measurements correspond to the spiral QST with $\{l=0\}$, followed by $\{l=0,1,-1\}$, $\{l=0,1,-1,2,-2\}$, and so on.
@@
In Fig. 4(a), we present the fidelity and the trace distance between the target state $\rho=|\Psi_{0}\rangle\langle\Psi_{0}|$ and the reconstructed state $\rho^{\rm rec}$ obtained via the spiral QST, with the number of repetitions (denoted by “reps”) fixed at ${\rm reps}=500$ for evaluating each expectation value. We compare the results for different amplitudes of magnetic field fluctuations ($\sigma_{\rm zp}=0.01,0.1,0.5$); other sources of noise are absent. First, we observe that the dependence on $\sigma_{\rm zp}$ appears to be absent, except for variations attributable only to statistical errors. Second, our “relevant-pitch-angles” strategy appears to be highly effective. The fidelity reaches approximately $0.98$ for $m/d^{2}\approx 0.1$, which corresponds to only 20-30 independent measurement setups, considering that one can obtain $d=2^{N}$ independent correlation functions from each experimental setup. The data points with the fewest measurements correspond to the results obtained by using just $m=3\times 2^{8}-2$ expectation values, based merely on three experimental setups of uniform $X$-, $Y$-, and $Z$-measurements (i.e., $l=0$ or $q=0$) in the SVT method. This condition already achieves very good fidelity and trace distance, likely due to the simplicity of the ground state of the NN Heisenberg model.
@@
In Fig. 4(b), we compare the results obtained by varying the number of repetitions for each experimental setting (i.e., the number of QGM shots after each unitary transformation corresponding to spiral operators defined by a specific pitch angle $q$ and spin plane) while keeping the amplitude of magnetic field fluctuations fixed at $\sigma_{\rm zp}=0.1$, although a different value would not change the results, as discussed above. We observe that this dependence is significant, at least below a certain threshold, although this sensitivity is a general issue for QST. Our results suggest that ${\rm reps}\gtrsim 500$ is required to achieve convergence.
@@
In Fig. 5, we present the results for the ground state of the $N=8$ Heisenberg + DM chain with $D=-J_{1}=1$ and $J_{p\neq 1}=0$, adopting the pitch-angle sequence $\{l=0\}$, $\{l=0,1\}$, $\{l=0,1,2\}$, etc. As seen in the figure, uniform measurements with $l=0$ alone are insufficient for accurately reconstructing the ground state in this case. The DM interaction introduces additional complexities, requiring the inclusion of finite pitch angles to capture the non-collinear spin correlations present in the ground state. Specifically, we identified the most relevant pitch angle (taken individually) in this case as $l=4$ or $q=\pi/2$ (see Appendix B).
[…]
III.2 Detection of entanglement measures
@@
for the reduced density matrix $\rho_{A}\equiv{\rm Tr}_{\bar{A}}(\rho)$ of subsystem $A$, when the whole system is in a pure quantum state. Here, ${\rm Tr}_{\bar{A}}(\cdot)$ denotes the partial trace with respect to the complement of $A$. The experimental detection of the entanglement entropy is of paramount importance, as it provides a versatile and precise measure of quantum correlations across various physical contexts. In condensed-matter physics, it is instrumental in identifying quantum critical points [53, 54] and topological order [55, 56]. In quantum field theory, it reveals underlying structural properties, including the central charge and operator content, through its relationship with the entanglement spectrum [57, 58]. Moreover, entanglement entropy is deeply connected to black hole entropy in gravitational theory [59, 60] and serves as a crucial indicator of information spread and thermalization processes in quantum thermodynamics [61, 62].
@@
Despite their critical importance, the experimental determination of entanglement measures remains challenging, particularly in large-scale systems. Direct measurements of entanglement have been rare, especially in more extensive systems. For instance, in cold-atom systems, while the second Rényi entanglement entropy $S_{2}(\rho_{A})$ has been measured by letting two identical copies of the quantum state in a Bose–Hubbard system interfere, such measurements have so far been implemented in systems with a modest number of sites [63, 61] (for an experimental measure of $S_{2}(\rho_{A})$ in trapped-ion systems see Ref. [64]).
@@
Figure 6: Spiral QST for the reduced density matrix. (a) Illustration of the bipartition of a system with NN sites into subsystem $A$, consisting of $N_{A}$ sites, and its complement $\bar{A}$. (b) Sample-averaged fidelity and trace distance for spiral compressed-sensing QST of the reduced density matrix $\rho_{A}={\rm Tr}_{\bar{A}}(\rho)$, where $N_{A}=7$, plotted as functions of the number of measurements $m$ scaled by $d^{2}=4^{7}$. The full density matrix $\rho$ corresponds to the ground state of the 14-site Heisenberg chain with NN antiferromagnetic spin-exchange coupling. The number of repetitions for evaluating each expectation value is set to ${\rm reps}=500$. The data points with the fewest measurements correspond to the spiral QST with $\{l=0\}$, followed by $\{l=0,1,-1\}$, $\{l=0,1,-1,2,-2\}$, and so on.
@@
We now demonstrate how entanglement measures in a quantum state can be extracted using spiral QST, which has the potential to overcome the challenge. Notably, to obtain the reduced density matrix $\rho_{A}$ via QST, it is not necessary to first reconstruct the entire pure state and then take a partial trace. Instead, the QST process can be applied directly to the measurement outcomes from the targeted subsystem $A$. To illustrate, we perform numerical simulations as follows: We consider a spin chain of length $N=14$, assumed to be prepared in the ground state $|\Psi_{0}\rangle$ of the Heisenberg model (8), discussed in the previous subsection. Focusing on a subsystem $A$ consisting of sites from $i=1$ to $i=N_{A}$ [see Fig. 6(a)], we simulate the reconstruction of the reduced density matrix $\rho_{A}^{\rm rec}$ on $A$ via the spiral compressed-sensing QST. The von Neumann and Rényi entanglement entropies are then calculated from the reconstructed $\rho_{A}^{\rm rec}$, and compared with their theoretical values.
@@
A priori, it is not clear whether spiral QST (or, more generally, compressed sensing) is suitable for a reduced density matrix, as it typically represents a mixed state of generic rank. Therefore, we begin by comparing $\rho_{A}^{\rm rec}$ with the theoretical reduced density matrix $\rho_{A}$, which is obtained by tracing out the complement of $A$ from $\rho=|\Psi_{0}\rangle\langle\Psi_{0}|$. In Fig. 6(b), we present the fidelity and trace distance between these matrices, with the subsystem size set to $N_{A}=7$ (half of the entire system). We set the number of repetitions for expectation value calculations to ${\rm reps}=500$. The value of $\sigma_{\rm zp}$ does not affect the results because the reduced density matrix retains the SU(2) symmetry of the ground state $|\Psi_{0}\rangle$ of the entire system. The results demonstrate that spiral QST is indeed effective also for the reduced density matrix, exhibiting a similar behavior to the case shown in Fig. 4. This can be attributed to the strong hierarchy of eigenvalues, where a few dominant ones significantly outweigh the others, a characteristic commonly observed in physical systems, which renders this reduced density matrix effectively low-rank.
@@
Now let us move on to the entanglement measures given in Eqs.(10) and (11). In Figs. 7(a-c), we present the sublattice-size $N_{A}$ dependence of $S_{\rm vN}(\rho_{A})$, $S_{2}(\rho_{A})$, and $S_{4}(\rho_{A})$, respectively, obtained from the reconstructed reduced density matrix $\rho_{A}^{\rm rec}$, for the case of ${\rm reps}=500$ (for concreteness $\sigma_{\rm zp}=0.1$ is chosen), alongside the theoretical values. Note that these are just representatives of the available entanglement measures; in fact, all the entanglement spectrum (and functions thereof) can be obtained from $\rho_{A}^{\rm rec}$ by simple diagonalization. The target state describes a subsystem of size $N_{A}$ of the 14-site NN antiferromagnetic Heisenberg chain ($J_{1}=1$ and $J_{p\neq 1}=0$) in its ground state. For the reconstruction, we used only the measurements with $l=0$ ($m=3\times 2^{N_{A}}-2$), as the fidelity and trace distance are already sufficiently high and low, respectively, as seen in Fig. 6(b).
@@
Moving beyond the previous simple model, we now demonstrate a more interesting case with antiferromagnetic NN coupling $J_{1}>0$ and next-nearest-neighbor (NNN) coupling $J_{2}>0$. These interactions compete with each other, leading to frustrations among the spins, which is expected to induce a larger entanglement in the ground state. In experiments with optical lattices, such a model can be realized by arranging the laser beams in a way that creates a zig-zag chain configuration [65]. Note that the zig-zag structure does not interfere with the procedure involving a magnetic field gradient along the chain direction.
We present the results for $J_{1}=J_{2}=1$ and $J_{p\neq 1,2}=0$ in Figs. 8(a-c). These plots reveal how well the spiral QST can capture the entanglement properties of the ground state of quantum many-body systems. Furthermore, the fact that the relative errors get better for higher Renyi index reinforces the notion that the reconstruction misses out only on the tail of small eigenvalues. These demonstrations provide valuable insights into the effectiveness of the method in important and interesting scenarios, such as frustrated quantum systems. To reinforce the efficiency of spiral QST in estimating this kind of nonlinear functions of the (reduced) density matrix, let us also note that the relative error achieved in Fig. 8, together with the low total number of measurements (including repetitions) used, is in line with the best results available in the literature for target states with similar amounts of entanglement (see in particular Fig. 4b in Ref. [28]).
IV Conclusions